OpenStudy (anonymous):

why total internal reflection happens only when a beam of light travels from a medium with higher refraction index to a medium with lower refraction index and not the other way around?

OpenStudy (siddharthbarua28):

The critical angle is the angle of incidence above which total internal reflection occurs. The angle of incidence is measured with respect to the normal at the refractive boundary (see diagram illustrating Snell's law). Consider a light ray passing from glass into air. The light emanating from the interface is bent towards the glass. When the incident angle is increased sufficiently, the transmitted angle (in air) reaches 90 degrees. It is at this point no light is transmitted into air. The critical angle \theta_c is given by Snell's law, n_1\sin\theta_i = n_2\sin\theta_t \quad. Rearranging Snell's Law, we get incidence \sin \theta_i = \frac{n_2}{n_1} \sin \theta_t. To find the critical angle, we find the value for \theta_i when \theta_t = 90° and thus \sin \theta_t = 1. The resulting value of \theta_i is equal to the critical angle \theta_c. Now, we can solve for \theta_i, and we get the equation for the critical angle: \theta_c = \theta_i = \arcsin \left( \frac{n_2}{n_1} \right), If the incident ray is precisely at the critical angle, the refracted ray is tangent to the boundary at the point of incidence. If for example, visible light were traveling through acrylic glass (with an index of refraction of approximately 1.50) into air (with an index of refraction of 1.00), the calculation would give the critical angle for light from acrylic into air, which is \theta _{c}=\arcsin \left( \frac{1.00}{1.50} \right)=41.8{}^\circ . Light incident on the border with an angle less than 41.8° would be partially transmitted, while light incident on the border at larger angles with respect to normal would be totally internally reflected. If the fraction {n_2}/{n_1} is greater than 1, then arcsine is not defined—meaning that total internal reflection does not occur even at very shallow or grazing incident angles. So the critical angle is only defined when {n_2}/{n_1} is less than 1.